OpenStudy (anonymous):

q = 868 − 20p Calculate the price elasticity of demand when the price is $31

OpenStudy (dumbcow):

\[E = \frac{p}{q}*\frac{dq}{dp}\] p=31 q = 868-620 = 248 dq/dp = -20